OpenStudy (anonymous):

Multiply the fraction and reduce to lowest terms: 16 x 5/9 x 2 1/2

OpenStudy (triciaal):

(16)(5)(21) _________ DIVIDE BY 2 AND DIVIDE BY 3 (9)(2) (8)(5)(7) / 3 =280/3 = 93 AND 1/3 WAS THAT SUPPOSE TO BE 21/2 OR 2 + 1/2 2+ 1/2 = 5/2 (16)(5)(5) / (9)(2)